OpenStudy (anonymous):

y= (sin4x) / (1+cos4x) Rewrite in the form of y=A TAN Bx y= A COT Bx y= A CSC Bx or y = A SEC Bx that has same graph.

OpenStudy (anonymous):

(sin4x) / (1+cos4x)= 2sin2x cos2x/2(cos 2x)^2 =tan2x so A

OpenStudy (anonymous):

so first one

OpenStudy (anonymous):

y = (sin 4x)/(1 + 4x) Replace sin 4x = 2*sin 2x*cos 2x (Identity: sin 2a = 2*sin a*cos a) Replace: 1 + cos 4x = 2*sin^2 2x (Identity: 1 + cos 2a = 2 sin^2 a) y = (2*sin 2x*cos 2a)/(2*sin^2 2x) = cos 2x/sin 2x = cot 2x

OpenStudy (anonymous):

Correction: (1 + cos 4a) = 2*cos ^2 2a y = sin 2x/cos 2x = tan 2x.